I measured the RC transmitted light signals here at the 40m. I made all connections through the PSL patch panel.
Other than two steering mirrors in front of the periscope, and the steering mirror for the RFPD which were used to steer
the beam into the cavity and the RFPD respectively, no optics are adjusted.
We re-aligned the beam into the cavity (the DC level increased from 2 V to 3.83V) (Fig2) (We could not recover the power back to what it was 90 days ago)
and the reflected beam to the center of the RFPD.
I measured the spectral density of the signal of the transmitted beam behind RefCav in both time and frequency domain.
This will be compared with the result from PSL lab later, so I can see how stable the signal should be.
I did not convert Vrms/rtHz to Hz/rtHz because I only look at the relative intensity of the transmitted beam which will be compared to the setup at PSL lab.
We care about this power fluctuation because we plan to measure
photo refractive noise on the cavity's mirros
(this is the noise caused by dn/dT in the coatings and the substrate,
the absorption from fluctuating power on the coating/mirror changes
the temperature which eventually changes the effective length of the cavity as seen by the laser.)
The plan is to modulate the power of the beam going into the cavity,
the absorption from ac part will induce frequency noise which we want to see.
Since the transmitted power of the cavity is proportional to the power inside the cavity.
Fluctuations from other factors, for example, gain setting, will limit our measurement.
That's why we are concerned about the stability of the transmitted beam and made this measurement.
|